# Geometric Hyperscanning of Affect under Active Inference

### v2

@@ 2.3 Geometric Hyperscanning as Second-Person Method
To empirically access the internal dynamics of dyadic affective inference, we introduce geometric hyperscanning based on the Forman-Ricci curvature (FRc) [23, 22], which captures phase transitions in inter-brain network topology using geometric markers derived from EEG hyperscanning data. This method enables operationalization of latent affective and narrative coherence as measurable transformations in the topology of inter-brain networks. Importantly, FRc is not treated as a post hoc readout; rather, we propose it as a data-derived proxy for latent generative states in the dyadic active inference model. Traditional hyperscanning methods focus on node-level synchrony, typically correlating homologous regions across two brains. While informative about temporal alignment, these methods fall short of detecting topological reconfigurations in inter-brain structure. FRc, by contrast, is an edge-centric geometric measure that quantifies the local expansion or contraction of topological flow within a network [45], and has been shown to characterize network robustness, signal routing, and functional reorganization.
@@
where we interpret $w(\cdot)$ as node and edge weights (e.g., derived from functional connectivity matrices). To track the evolution of inter-brain topology over time, we calculate the entropy of the FRc distribution within a sliding window.
@@
FRc entropy is not simply an output variable but can be formalized as an observation model over latent affective prediction error. That is, curvature dynamics can be used to infer or constrain internal state variables such as valence or identity coherence. This closes the loop between model and measurement, as the generative model predicts changes in self-model coherence and dyadic free energy, and the FRc entropy serves as an empirical signal that can be used to update model beliefs via Bayesian inference. To make these transitions tractable and interpretable, we propose a multi-level operationalization of rupture, repair, and reattunement, linking neural signatures to behavioral and psychological indicators. Neurally, rupture corresponds to abrupt drops in network integration (low FRc), high curvature entropy, or rapid shifts in entropy gradients; repair and reattunement reflect a return to more stable topologies. Behaviorally, rupture may manifest as disengagement, disrupted synchrony, or confusion (e.g., gaze aversion, vocal interruptions), while repair involves re-engagement cues such as gesture mirroring, prosodic modulation, or turn-taking; reattunement is evidenced by fluent, coordinated interaction and mutual responsiveness. Psychologically, rupture can be self-reported as disconnection or emotional distancing, repair as subjective efforts to reconnect, and reattunement as restored mutual understanding or relational safety. These experiences can be measured via self-reports (affective sliders, post-session interviews, validated instruments), annotated through multimodal behavioral data (gaze, prosody, synchrony), and aligned with curvature dynamics in the inter-brain manifold. Treated as latent categorical states (e.g., within a hidden Markov model), these phases can serve as points of annotation, classification, or real-time feedback in both experimental and clinical contexts.
@@
Importantly, embedding these events within the recursive generative model of dyadic inference allows rupture–repair cycles to function not just as outcomes but as endogenous features of interactional dynamics. Affective coherence drives policy selection, while behavioral and neural feedback recursively update each agent’s self-model. This recursive structure enables in situ inference over dynamic interpersonal states, validation of model predictions across modalities, and fine-grained adaptation to moment-by-moment shifts in relational coherence.Integrating geometric hyperscanning into the generative model therefore bounds the model with an additional objective: not only minimising identity-relevant prediction error, but implicitly driving the system toward maximal interpersonal attunement—an intrinsically rewarding state even in the absence of overt instrumental gain [35, 7]. Curvature-induced surprises can surface to awareness, prompting explicit attempts at repair.
@@
In sum, FRc-based geometric hyperscanning offers a scalable, multimodal method for detecting and interpreting recursive affective inference. By triangulating neural entropy, behavioral dynamics, and psychological experience, we gain an integrated empirical handle on how dyads co-regulate meaning, affect, and identity over time. This framework not only advances a formal, testable architecture for second-person neuroscience but also lays the groundwork for modeling recursive emotional entanglement, the subject of the following section.
@@ 3 Future Directions and Conclusions
In this paper, we proposed that affect is best understood as recursive inference over self-model coherence within coupled generative systems. Modeling the dyad via active inference, we formalized affect as a valuation of identity-relevant surprise, driving and modulating recursive loops of belief updating, policy selection, and behavioral adaptation during social interaction. We introduced a novel method based on the FRc, termed geometric hyperscanning [22], which tracks dynamic reconfigurations in inter-brain networks, providing an empirical window onto moments of rupture, repair, and re-attunement. By unifying formal models of belief dynamics, affective evaluation, and network geometry, we advance a scalable and interpretable framework for operationalizing second-person active inference.
@@ 3.1 Future Directions
The theoretical formulation proposed here invites a number of concrete avenues for empirical investigation and computational development. Foremost among these is the simulation of agent-based models in which curvature entropy is treated as a streamed sensory input—continuously informing belief updates about narrative alignment. Such models would allow us to test whether the architecture, under active inference, spontaneously reproduces the rupture–repair cycles observed in real dyadic interaction. In doing so, they provide a crucial bridge between formalism, neurophysiology, and phenomenology. Beyond dyads, the framework naturally extends to hierarchically structured systems. By introducing group-level alignment variables one can model small-group coherence as an emergent property of multiple interacting pairwise dynamics. This multiscale extension opens the door to applications in collective behaviour, organisational synchrony, and crowd psychology, where relational inference is distributed across agents and temporal scales. Furthermore, the curvature-informed likelihood may be productively integrated with other sensorimotor channels. Prosodic variation, gaze dynamics, and even facial microexpressions can all be incorporated within a deep generative model, allowing heterogeneous evidence streams to converge on shared latent variables. This multimodal fusion enriches the inferential landscape and supports more robust decoding of the intersubjective field. Yet the prospect of real-time inference over relational states is not without ethical consequence. As structural priors become actionable in applied contexts—be it psychotherapy, education, or social robotics—the issues of privacy, autonomy, and interpretability move to the fore. Any translational implementation must therefore be grounded in transparent consent procedures, and designed to yield feedback that is intelligible not only to clinicians and researchers, but also to the individuals whose relational trajectories it purports to track.
@@ 3.2 Sociomarkers for Interpersonalized Psychiatry
Geometric hyperscanning offers a promising foundation for sociomarkers: real-time, measurable signatures of dynamic interpersonal coordination [6], analogous to biomarkers of individual physiology. Affective co-regulation—vital for social bonding across species [12, 16, 8, 36]—is foundational to human development, shaping cognition, emotion, and relational selfhood through early caregiver-infant interactions [7, 15, 43, 17, 24]. Misattunement and repair cycles scaffold resilience and relational growth [6, 7], with fluctuations in curvature entropy potentially indexing key developmental processes [37, 11, 44]. Such signatures could inform personalized [31, 30] and interpersonalized psychiatry [6], which views psychopathology as disruption in co-regulation and narrative repair [5, 32]. In psychotherapy, real-time dyadic metrics [5, 28] can track attunement, rupture, and repair [34], helping both therapists and clients navigate the relational dynamics of change. Crucially, synchrony is neither inherently good nor constant: transitions into and out of synchrony may reflect autonomy, resistance, or meaningful differentiation, not dysfunction. Modeling these dynamics as recursive affective inference provides a principled lens for both basic science and relational mental health care.
@@ 3.3 Conclusions
Building on these formal and empirical insights, this approach carries several key implications. First, it invites a rethinking of affect in social interaction—not merely as a byproduct of inference, but as a primary regulatory signal that modulates the stability and flexibility of generative coupling itself. On this view, affect becomes the felt sense of coherence or incoherence within the dyadic narrative, serving as a signal for when beliefs must be renegotiated or when interactional roles require repair. Second, by shifting emphasis from synchrony per se to the topology of inter-brain networks, our method aligns with a broader movement in second-person neuroscience toward characterizing the structure of interaction. In this respect, FRc extends Friston and Frith’s (2015) [18] notion of generalized synchrony by capturing the reconfigurations in network structure that accompany recursive affective inference. Third, by embedding affect into the posterior over policies, we extend da Costa’s (2020) planning-as-inference formalism into the affective domain. Crucially, the affective signal can operate implicitly, as a precision-weight on policy beliefs that never reaches phenomenal awareness, or explicitly, as a consciously accessible feeling that furnishes the agent with propositional knowledge about its own relational stance. In either guise, policy selection is driven not only by epistemic or instrumental imperatives, but also by the imperative to preserve narrative and relational coherence. Affect thereby becomes an intrinsic dimension of social planning, covertly biasing action when tacit and overtly guiding behaviour when made conscious—steering agents through moments of ambiguity toward restored interactional alignment. These considerations open a range of empirical and computational possibilities. In psychotherapy, curvature entropy may serve as a real-time sociomarker of rupture and repair; in early development, FRc could illuminate how co-regulatory patterns emerge between infants and caregivers. Simulations of affective agents equipped with recursive self- and other-models offer a formal means to investigate relational resilience and breakdown. In particular, such models allow us to test the dialectical misattunement hypothesis (Bolis et al., 2017), which posits that psychiatric risk arises not from individual deficits alone but from persistent mismatches in communicative expectations or learning rates. Simulating these mismatches under controlled conditions reveals how dyads adapt—or fail to—over developmental time, providing insight into both vulnerability and repair mechanisms. The theoretical landscape also invites expansion. How does this framework scale to group interactions, where recursive generative coupling involves multiple agents? What ethical safeguards are needed when automating the detection of affective rupture and repair—especially in sensitive contexts like therapy or education? And how might shared belief geometry interact with linguistic, prosodic, and bodily cues to form a rich, multimodal map of affective inference? Together, these questions point toward a broader scientific project: one that integrates formal modeling, empirical measurement, and philosophical reflection into a unified account of relational meaning. In this view, the dynamics of belief, affect, and participation are not peripheral but fundamental to the fabric of social cognition. By reframing affect as recursive inference within coupled generative systems—and rendering those dynamics tractable via network geometry—we move closer to the vision of interaction as the basic unit of social neuroscience. Our proposal thus offers both a conceptual framework and methodological toolkit for studying the structural and affective choreography of human interaction.

### v3

@@ 2.3 Geometric Hyperscanning as Second-Person Method
To empirically access the internal dynamics of dyadic affective inference, we introduce geometric hyperscanning based on the Forman-Ricci curvature (FRc) [22, 23], which captures phase transitions in inter-brain network topology using geometric markers derived from EEG hyperscanning data. This method enables operationalization of latent affective and narrative coherence as measurable transformations in the topology of inter-brain networks. Importantly, FRc is not treated as a post hoc readout; we propose it as a data-derived proxy for latent generative states in the dyadic active inference model. Traditional hyperscanning methods focus on node-level synchrony, typically correlating homologous regions across two brains. While informative about temporal alignment, these methods fail to detect topological reconfigurations in inter-brain structure. FRc, by contrast, is an edge-centric geometric measure that quantifies the local expansion or contraction of topological flow within a network [45], and has been shown to characterize network robustness, signal routing, and functional reorganization.
@@
where we interpret $w(\cdot)$ as node and edge weights (e.g., derived from functional connectivity matrices), to track the evolution of inter-brain topology over time, we calculate the entropy of the FRc distribution within a sliding window.
@@
FRc entropy is not simply an output variable but can be formalized as an observation model over latent affective prediction error. Curvature dynamics can infer or constrain internal state variables such as valence or identity coherence. This closes the loop between model and measurement, as the generative model predicts changes in self-model coherence and dyadic free energy, and the FRc entropy serves as an empirical signal that can be used to update model beliefs via Bayesian inference. To make these transitions tractable and interpretable, we propose a multi-level operationalization of rupture, repair, and reattunement, linking neural signatures to behavioral and psychological indicators. Neurally, rupture corresponds to abrupt drops in network integration (low FRc), high curvature entropy, or rapid shifts in entropy gradients; repair and reattunement reflect a return to more stable topologies. Behaviorally, rupture may manifest as disengagement, disrupted synchrony, or confusion (e.g., gaze aversion, vocal interruptions), while repair involves re-engagement cues such as gesture mirroring, prosodic modulation, or turn-taking; reattunement is evidenced by fluent, coordinated interaction and mutual responsiveness. Psychologically, rupture can be self-reported as disconnection or emotional distancing, repair as subjective efforts to reconnect, and reattunement as restored mutual understanding or relational safety. These experiences can be measured via self-reports (affective sliders, post-session interviews, validated instruments), annotated through multimodal behavioral data (gaze, prosody, synchrony), and aligned with curvature dynamics in the inter-brain manifold. Treated as latent categorical states (e.g., within a hidden Markov model), these phases can serve as points of annotation, classification, or real-time feedback in both experimental and clinical contexts.
@@
Importantly, embedding these events within the recursive generative model of dyadic inference allows rupture–repair cycles to function as outcomes and as endogenous features of interactional dynamics. Affective coherence drives policy selection, while behavioral and neural feedback recursively update each agent’s self-model. This recursive structure enables in situ inference over dynamic interpersonal states, validation of model predictions across modalities, and fine-grained adaptation to moment-by-moment shifts in relational coherence. Integrating geometric hyperscanning into the generative model therefore bounds the model with an additional objective: not only minimising identity-relevant prediction error, but implicitly driving the system toward maximal interpersonal attunement—an intrinsically rewarding state even in the absence of overt instrumental gain [35, 7]. Curvature-induced surprises can surface to awareness, prompting explicit attempts at repair.
@@
In sum, FRc-based geometric hyperscanning offers a scalable, multimodal method for detecting and interpreting recursive affective inference. By triangulating neural entropy, behavioral dynamics, and psychological experience, we gain an integrated empirical handle on how dyads co-regulate meaning, affect, and identity over time. This framework advances a formal, testable architecture for second-person neuroscience and lays the groundwork for modeling recursive emotional entanglement, the subject of the following section.
@@ 3 Future Directions and Conclusions
This paper proposed that affect is best understood as recursive inference over self-model coherence within coupled generative systems. Modeling the dyad via active inference, we formalized affect as a valuation of identity-relevant surprise, driving and modulating recursive loops of belief updating, policy selection, and behavioral adaptation during social interaction. We introduced a novel method based on the FRc, termed geometric hyperscanning [23], which tracks dynamic reconfigurations in inter-brain networks, providing an empirical window onto moments of rupture, repair, and re-attunement. By unifying formal models of belief dynamics, affective evaluation, and network geometry, we advance a scalable and interpretable framework for operationalizing second-person active inference.
@@ 3.1 Future Directions
The proposed theoretical formulation invites several concrete avenues for empirical investigation and computational development. Foremost among these is the simulation of agent-based models in which curvature entropy is treated as a streamed sensory input—continuously informing belief updates about narrative alignment. Such models would allow us to test whether the architecture, under active inference, spontaneously reproduces the rupture–repair cycles observed in real dyadic interaction. In doing so, they provide a crucial bridge between formalism, neurophysiology, and phenomenology. Beyond dyads, the framework naturally extends to hierarchically structured systems. By introducing group-level alignment variables, one can model small-group coherence as an emergent property of multiple interacting pairwise dynamics. This multiscale extension opens the door to collective behaviour, organisational synchrony, and crowd psychology applications, where relational inference is distributed across agents and temporal scales. Furthermore, the curvature-informed likelihood may be productively integrated with other sensorimotor channels. Prosodic variation, gaze dynamics, and even facial microexpressions can all be incorporated within a deep generative model, allowing heterogeneous evidence streams to converge on shared latent variables. This multimodal fusion enriches the inferential landscape and supports more robust decoding of the intersubjective field. Yet the prospect of real-time inference over relational states is not without ethical consequences. As structural priors become actionable in applied contexts—be it psychotherapy, education, or social robotics—the issues of privacy, autonomy, and interpretability move to the fore. Any translational implementation must therefore be grounded in transparent consent procedures and designed to yield intelligible feedback to clinicians and researchers and to the individuals whose relational trajectories it purports to track.
@@ 3.2 Sociomarkers for Interpersonalized Psychiatry
Geometric hyperscanning offers a promising foundation for sociomarkers: real-time, measurable signatures of dynamic interpersonal coordination [6], analogous to biomarkers of individual physiology. Affective co-regulation—vital for social bonding across species [12, 16, 8, 36]—is foundational to human development, shaping cognition, emotion, and relational selfhood through early caregiver-infant interactions [7, 15, 43, 17, 24]. Misattunement and repair cycles scaffold resilience and relational growth [6, 7], with fluctuations in curvature entropy potentially indexing key developmental processes [37, 11, 44]. Such signatures could inform personalized [31, 30] and interpersonalized psychiatry [6], which views psychopathology as a disruption in co-regulation and narrative repair [5, 32]. In psychotherapy, real-time dyadic metrics [5, 28] can track attunement, rupture, and repair [34], helping both therapists and clients navigate the relational dynamics of change. Crucially, synchrony is neither inherently good nor constant: transitions into and out of synchrony may reflect autonomy, resistance, or meaningful differentiation, not dysfunction. Modeling these dynamics as recursive affective inference provides a principled lens for basic science and relational mental health care.
@@ 3.3 Conclusions
Building on these formal and empirical insights, this approach carries several key implications. First, it invites a rethinking of affect in social interaction—not merely as a byproduct of inference, but as a primary regulatory signal that modulates the stability and flexibility of generative coupling itself. On this view, affect becomes the sense of coherence or incoherence within the dyadic narrative, indicating when beliefs must be renegotiated or when interactional roles require repair. Second, by shifting emphasis from synchrony per se to the topology of inter-brain networks, our method aligns with a broader movement in second-person neuroscience toward characterizing interaction structure. In this respect, FRc extends Friston and Frith’s (2015) [18] notion of generalized synchrony by capturing the reconfigurations in network structure that accompany recursive affective inference. Third, by embedding affect into the posterior over policies, we extend da Costa’s (2020) planning-as-inference formalism into the affective domain. Crucially, the affective signal can operate implicitly, as a precision-weight on policy beliefs that never reaches phenomenal awareness, or explicitly, as a consciously accessible feeling that furnishes the agent with propositional knowledge about its own relational stance. In either guise, policy selection is driven not only by epistemic or instrumental imperatives, but also by the imperative to preserve narrative and relational coherence. Affect thereby becomes an intrinsic dimension of social planning, covertly biasing action when tacit and overtly guiding behaviour when made conscious—steering agents through moments of ambiguity toward restored interactional alignment. These considerations open a range of empirical and computational possibilities. In psychotherapy, curvature entropy may serve as a real-time sociomarker of rupture and repair; in early development, FRc could illuminate how co-regulatory patterns emerge between infants and caregivers. Simulations of affective agents equipped with recursive self- and other-models offer a formal means to investigate relational resilience and breakdown. In particular, such models allow us to test the dialectical misattunement hypothesis (Bolis et al., 2017), which posits that psychiatric risk arises not from individual deficits alone but from persistent mismatches in communicative expectations or learning rates. Simulating these mismatches under controlled conditions reveals how dyads adapt—or fail to—over developmental time, providing insight into vulnerability and repair mechanisms. The theoretical landscape also invites expansion. How does this framework scale to group interactions, where recursive generative coupling involves multiple agents? What ethical safeguards are needed when automating the detection of affective rupture and repair—especially in sensitive contexts like therapy or education? And how might shared belief geometry interact with linguistic, prosodic, and bodily cues to form a rich, multimodal map of affective inference? Together, these questions point toward a broader scientific project: one that integrates formal modeling, empirical measurement, and philosophical reflection into a unified account of relational meaning. In this view, the dynamics of belief, affect, and participation are not peripheral but fundamental to the fabric of social cognition. By reframing affect as recursive inference within coupled generative systems—and rendering those dynamics tractable via network geometry—we move closer to the vision of interaction as the basic unit of social neuroscience. Our proposal thus offers both a conceptual framework and methodological toolkit for studying human interaction’s structural and affective choreography.